### v2

# Mott-glass phase induced by long-range correlated disorder in a one-dimensional Bose gas

Abstract
We determine the phase diagram of a one-dimensional Bose gas in the presence of disorder with short- and long-range correlations, the latter decaying with distance as $1/|x|^{1+\sigma}$. When $\sigma<0$, the Berezinskii-Kosterlitz-Thouless transition between the superfluid and the localized phase is driven by the long-range correlations and the Luttinger parameter $K$ takes the critical value $K_{c}(\sigma)=3/2-\sigma/2$. The localized phase is a Bose glass for $\sigma>\sigma_{c}=3-\pi^{2}/3\simeq-0.289868$, and a Mott glass —characterized by a vanishing compressibility and a gapless conductivity— when $\sigma<\sigma_{c}$. Our conclusions, based on the nonperturbative functional renormalization group and perturbative renormalization group, are confirmed by the study of the case $\sigma=-1$, corresponding to a perfectly correlated disorder in space, where the model is exactly solvable in the semiclassical limit $K\to 0^{+}$.
Introduction.
Long-range interactions, decaying as a power law with distance, have crucial consequences on the physical properties of condensed-matter systems. They can alter the phase diagram, the critical properties of phase transitions and the dynamical behavior of physical observables [1]. In disordered systems, long-range correlations of the defects play a similar role since they lead, after averaging over the disorder, to effective long-range (elastic) interactions. The modification of the critical behavior and the stabilization of new phases by a long-range correlated disorder has been studied in many systems, including spin systems with random-bond [2, 3, 4, 5, 6, 7] and random-field [8, 9] disorder, different Anderson insulators [10, 11, 12], topological nodal semimetals [13, 14, 15], and random elastic manifolds such as domain walls in ferromagnets and flux lines in superconductors [16, 17, 18, 19, 20, 21].
Figure 1: Phase diagram of a one-dimensional Bose gas in the presence of a (weak) disorder with long-range correlations decaying with distance as $1/|x|^{1+\sigma}$. The critical value separating the Bose glass from the Mott glass is $\sigma_{c}=3-\pi^{2}/3\simeq-0.289868$. The gray area shows the region where the FRG approach [Eqs. (9-10)] breaks down. The thick line at $\sigma=-1$ corresponds to the case of a perfectly correlated disorder, where the exact solution in the semiclassical limit $K\to 0^{+}$ confirms that the ground state is a Mott glass
In this Letter, we study the phase diagram of a one-dimensional Bose gas in the presence of short- and long-range disorder correlations, the latter decaying with distance as $1/|x|^{1+\sigma}$, using bosonization, the replica formalism and the nonperturbative functional renormalization group (FRG) (similar results are obtained with the perturbative field-theoretical FRG [22]). Our main results are summarized in Fig. 1. The Berezinskii-Kosterlitz-Thouless transition between the superfluid and the localized phase, which occurs for $K_{c}=3/2$ (for weak disorder) in the absence of long-range correlations, occurs for $K_{c}(\sigma)=3/2-\sigma/2$ when the exponent $\sigma$ is negative. The long-range disorder correlations are nevertheless irrelevant in the localized phase when $\sigma>\sigma_{c}=3-\pi^{2}/3\simeq-0.289868$ and the system is a Bose glass with a nonzero compressibility and a gapless conductivity [23, 24]. The “short-range” fixed point is unstable when $\sigma<\sigma_{c}$. We find that the localized phase is described by a “long-range” fixed point, corresponding to a Mott glass with a vanishing compressibility and a gapless conductivity, but point out that the cumulant expansion used in the RG approach formally breaks down. In the case $\sigma=-1$, corresponding to a perfectly correlated disorder in space, the model in exactly solvable in the semiclassical limit $K\to 0^{+}$ and we find a vanishing compressibility and a gapless conductivity varying at $|\omega|^{3}$ in the small-frequency limit. While this result strongly supports the existence of a Mott glass for $\sigma<\sigma_{c}$, we cannot exclude a quantum phase transition at a critical value $-1\leq\sigma^{*}<\sigma_{c}$ between two incompressible localized phases with different low-frequency conductivity (e.g. $\omega^{2}$ vs $|\omega|^{3}$).
Model and FRG formalism.
We consider a one-dimensional Bose gas described by the Luttinger-liquid Hamiltonian
[…]
in the bosonization formalism [25], where $K$ is the Luttinger parameter and $v$ the velocity of the sound mode. The operator $\hat{\theta}$ is the phase of the boson operator $\hat{\psi}=e^{i\theta}\sqrt{\hat{\rho}}$ and $\partial_{x}\hat{\varphi}$ is conjugate to $\hat{\theta}(x)$: $[\hat{\theta}(x),\partial_{y}\hat{\varphi}(y)]=i\pi\delta(x-y)$. A random potential that couples to the boson density is taken into account by the Hamiltonian
$\hat{H}_{\rm dis}=\int dx\left[-\frac{1}{\pi}\eta\partial_{x}\varphi+(\rho_{2}\xi^{*}e^{2i\hat{\varphi}}+{\rm H.c.})\right],$ (2)
where $\eta(x)$ (real) and $\xi(x)$ (complex) denote random potentials with Fourier components near 0 and ±$\pm 2\pi\rho_{0}$, respectively [26, 27]. $\rho_{0}$ is the average particle density and $\rho_{2}$ a nonuniversal parameter that depends on microscopic details. $\eta(x)$ can be eliminated by a shift of $\varphi$ and is not considered in the following. The random potential $\xi(x)$ is assumed to be Gaussian correlated with zero mean and variance
$\begin{gathered}\overline{\xi(x)\xi(x^{\prime})}=\overline{\xi^{*}(x)\xi^{*}(x^{\prime})}=0,\\ \overline{\xi^{*}(x)\xi(x^{\prime})}=D_{\rm SR}\delta(x-x^{\prime})+D_{\rm LR}f(x-x^{\prime}),\end{gathered}$ (3)
where the overbar denotes the average over disorder. The delta correlated part corresponds to short-range correlations while $f(x)\sim 1/|x|^{1+\sigma}$ describes long-range correlations. When $\sigma>0$, the Fourier transform $f(q)$ of the function $f(x)$ has a well-defined limit when $q\to 0$ and $f(x)$ is effectively short-range, merely renormalizing the short-range part described by $D_{\rm SR}$. In the following we assume $\sigma\leq 0$ so that $f(q)$ diverges when $q\to 0$. We normalize $f(x)$ so that $f(q)=|q|^{\sigma}$ for $q\to 0$ if $\sigma<0$ and $f(q\to 0)=\ln(1/|q|)$ if $\sigma=0$. One can set $D_{\rm SR}=0$ if the disorder has no short-range correlations, however a nonzero $D_{\rm SR}$ is always generated under coarse graining [17].
We average over disorder by introducing $n$ replicas of the system [27, 24, 28]. After integrating out the phase field $\theta$, we obtain the Euclidean action
$S[\{\varphi_{a}\}]=\sum_{a}\int_{x,\tau}\frac{v}{2\pi K}\left[(\partial_{x}\varphi_{a})^{2}+\frac{({\partial_{\tau}}\varphi_{a})^{2}}{v^{2}}\right]\\ -\sum_{a,b}\int_{x,x^{\prime},\tau,\tau^{\prime}}[{\cal D}_{\rm SR}\delta(x-x^{\prime})\\ +{\cal D}_{\rm LR}f(x-x^{\prime})]\cos[2\varphi_{a}(x,\tau)-2\varphi_{b}(x^{\prime},\tau^{\prime})],$ (4)
[…]
Refer to caption Refer to caption
Figure 2: Flow diagrams in the space $(K,\sqrt{\lambda_{2}},\sqrt{\delta_{1}})$ for $\sigma=-0.2$ (left) and $\sigma=-0.4$ (right). The black dots show the BKT fixed point $((3-\sigma)/2,0,0)$ controlling the transition between the superfluid phase and the localized phase, as well as the Bose-glass (BG) and Mott-glass (MG) fixed points. (The dashing of some lines and the different colors are used to improve visibility and have no particular physical meaning.)
Most physical quantities can be obtained from the effective action (or Gibbs free energy),
$\Gamma[\{\phi_{a}\}]=-\ln Z[\{J_{a}\}]+\sum_{a}\int_{x,\tau}J_{a}\phi_{a},$ (5)
defined as the Legendre transform of the free energy $-\ln Z[\{J_{a}\}]$. Here $J_{a}$ is an external source which couples linearly to the field $\varphi_{a}$ and allows us to obtain the expectation value $\phi_{a}(x,\tau)=\langle\varphi_{a}(x,\tau)\rangle=\delta\ln Z[\{J_{a}\}]/\delta J_{a}(x,\tau)$. We compute $\Gamma[\{\phi_{a}\}]$ using a nonperturbative FRG approach [29, 30, 31]. In practice, we consider a scale-dependent effective action $\Gamma_{k}[\{\phi_{a}\}]$ which includes fluctuations with momenta (and frequencies) between the UV cutoff $\Lambda$ of the theory and a running momentum scale $k\leq\Lambda$. The effective action $\Gamma_{k=0}[\{\phi_{a}\}]$ of the original model is obtained when all fluctuations have been integrated out whereas $\Gamma_{\Lambda}[\{\phi_{a}\}]=S[\{\phi_{a}\}]$ is given by the microscopic action. $\Gamma_{k}$ satisfies a flow equation which allows us to obtain $\Gamma_{k=0}$ from $\Gamma_{\Lambda}$ but which cannot be solved exactly [32, 33, 34]. We refer to the Supplemental Material [22] for a more detailed presentation of the FRG approach as well as a discussion of the perturbative field-theoretical approach.
We consider the following truncation of the effective action,
$\Gamma_{k}[\{\phi_{a}\}]=\sum_{a}\Gamma_{1,k}[\phi_{a}]-\frac{1}{2}\sum_{a,b}\Gamma_{2,k}[\phi_{a},\phi_{b}],$ (6)
with the ansatz
$\Gamma_{1,k}[\phi_{a}]=\int_{x,\tau}\frac{v_{k}}{2\pi K_{k}}\left[(\partial_{x}\phi_{a})^{2}+\frac{({\partial_{\tau}}\phi_{a})^{2}}{v_{k}^{2}}\right]$ (7)
and
$\mskip-3.0mu\mskip-3.0mu\mskip-3.0mu\mskip-3.0mu\mskip-3.0mu\Gamma_{2,k}[\phi_{a},\phi_{b}]=\int_{x,x^{\prime},\tau,\tau^{\prime}}\bigl{[}\delta(x-x^{\prime})V_{1,k}(\phi_{a}(x,\tau)-\phi_{b}(x^{\prime},\tau^{\prime}))\\ +f(x-x^{\prime})V_{2,k}(\phi_{a}(x,\tau)-\phi_{b}(x^{\prime},\tau^{\prime}))\bigr{]}.$ (8)
Equation (6) is based on a free-replica-sum expansion where all “cumulants” $\Gamma_{i,k}$ with $i\geq 3$ are ignored. The form of $\Gamma_{1,k}$ follows from a derivative expansion to second order. The coefficients of the derivative terms are written in terms of a Luttinger parameter $K_{k}$ and a velocity $v_{k}$ which are scale dependent. This defines a scale-dependent superfluid stiffness $\rho_{s,k}=v_{k}K_{k}/\pi$ and a scale-dependent compressibility $\kappa_{k}=K_{k}/\pi v_{k}$. The two-replica part $\Gamma_{2,k}$ is assumed to keep the same form as the two-replica term in the microscopic action but with renormalized potentials $V_{1,k}$ and $V_{2,k}$. The initial conditions are $K_{\Lambda}=K$, $v_{\Lambda}=v$, $V_{1,\Lambda}(u)=2{\cal D}_{\rm SR}\cos(2u)$ and $V_{2,\Lambda}(u)=2{\cal D}_{\rm LR}\cos(2u)$.
The flow equations are obtained by inserting the ansatz (6-8) into the exact flow equation satisfied by the effective action,
$k\partial_{k}\Delta_{1}={}$ $(-3+2\theta-2z+2)\Delta_{1}-Kl_{1}\Delta_{1}^{\prime\prime}$
$+\bar{L}_{0}[\Delta_{1}^{\prime}{}^{2}+\Delta_{1}^{\prime\prime}(\Delta_{1}-\Delta(0))]$
$+\bar{L}_{1}[2\Delta_{1}^{\prime}\Delta_{2}^{\prime}+\Delta_{1}^{\prime\prime}(\Delta_{2}-\Delta_{2}(0))+\Delta_{1}\Delta_{2}^{\prime\prime}]$
$+\bar{L}_{2}(\Delta_{2}^{\prime}{}^{2}+\Delta_{2}\Delta_{2}^{\prime\prime}),$ (9)
$k\partial_{k}\Delta_{2}={}$ $(-3+\sigma+2\theta-2z+2)\Delta_{2}-Kl_{1}\Delta_{2}^{\prime\prime}$
$-\bar{L}_{0}\Delta_{2}^{\prime\prime}\Delta_{1}(0)-\bar{L}_{1}\Delta_{2}^{\prime\prime}\Delta_{2}(0)$ (10)
(we have dropped the $k$ index in $\Delta_{i}$, $K$, $\theta$, and $z$ for ease of notations), where
$\begin{split}\Delta_{1,k}(u)&=-\frac{K_{k}^{2}V^{\prime\prime}_{1,k}(u)}{v_{k}^{2}k^{3}},\;\;\;\Delta_{2,k}(u)=-\frac{K_{k}^{2}V^{\prime\prime}_{2,k}(u)}{v_{k}^{2}k^{3-\sigma}},\end{split}$ (11)
are dimensionless potentials, and $\theta_{k}=k\partial_{k}\ln K_{k}$, $z_{k}-1=k\partial_{k}\ln v_{k}$ ($z_{k}$ is the dynamical exponent),
$\theta_{k}$ $=\frac{1}{2}[\bar{M}_{\tau}\Delta_{1,k}^{\prime\prime}(0)+(\bar{L}_{x}+\bar{L}_{\tau})\Delta_{2,k}^{\prime\prime}(0)],$ (12)
$z_{k}-1$ $=\frac{1}{2}[\bar{M}_{\tau}\Delta_{1,k}^{\prime\prime}(0)-(\bar{L}_{x}-\bar{L}_{\tau})\Delta_{2,k}^{\prime\prime}(0)].$ (13)
The “threshold functions” $\bar{L}_{i}$ ($i\in[1,3]$), $\bar{L}_{x}$, $\bar{L}_{\tau}$ and $\bar{M}_{\tau}$ are $k$-independent linear functions of $\theta_{k}$ and $z_{k}$ [22]. In the absence of long-range correlated disorder, the coefficient $v_{k}/2\pi K_{k}$ of the space derivative term in $\Gamma_{1}$ does not renormalize, in agreement with the statistical tilt symmetry of the action (4[24, 35].
Equations (9-13) can be solved numerically expanding the $\pi$-periodic functions $\Delta_{1,k}(u)=\sum_{n=1}^{\infty}\delta_{n,k}\cos(2nu)$ and $\Delta_{2,k}(u)=\lambda_{2,k}\cos(2u)$ in circular harmonics (with an upper cutoff $n_{\rm max}<\infty$ on the number harmonics). The function $\Delta_{2,k}(u)$ has a single harmonics when $k=\Lambda$ and this form is obviously preserved by the flow equation (10). Figure 2 shows RG flows, for $\sigma=-0.2$ and $\sigma=-0.4$, projected onto the plane $(K,\sqrt{\lambda_{2}},\sqrt{\delta_{1}})$. In both cases one observes a critical BKT fixed point controlling a transition between a superfluid phase and a localized phase. For $\sigma=-0.2$, the localized phase is described by an attractive fixed point where the long-range correlated disorder vanishes, $\lambda_{2}^{*}=0$, and corresponds to the Bose-glass phase described in previous studies [23, 36]. For $\sigma=-0.4$, the long-range part of the disorder does not vanish at the fixed point describing the localized phase ($\lambda_{2}^{*}\neq 0$); we shall see that this fixed point describes a Mott glass.
Superfluid-insulator transition.
The stability of the superfluid phase against an infinitesimal disorder can be determined from scaling analysis considering the microscopic action (4). At the Luttinger-liquid fixed point, $e^{2i\varphi}$ has scaling dimension $K$, so that $[D_{\rm SR}]=3-2K$ and $[D_{\rm LR}]=3-\sigma-2K$ using $[x]=[\tau]=-1$ and $[f(x)]=1+\sigma$. The long-range correlated part of the disorder becomes relevant for $K<K_{c}(\sigma)=3/2-\sigma/2$ and $\sigma<0$, destabilizing the superfluid phase, so that the critical value of the Luttinger parameter, $K_{c}(\sigma)$, is larger than the value $3/2$ obtained for disorder with only short-range correlations. By neglecting the short-range correlated disorder, which is irrelevant near the BKT fixed point when $\sigma<0$, the flow equations (9-13) reproduce the standard RG equations of the BKT theory to leading order in $K_{k}-K_{c}(\sigma)$ and $\sqrt{\lambda_{2,k}}$ [22]. These equations are in very good agreement with the numerical solution of the full equations (Fig. 3).
Figure 3: Flow trajectories, projected in the plane $(K,\sqrt{\lambda_{2}})$, near the BKT fixed point located at $K_{c}=1.7$ ($\sigma=-0.4$). The initial conditions are defined by $\delta_{1,\Lambda}\sim 0.1$ when $K_{\Lambda}>1.7$, and $\delta_{1,\Lambda}\sim 10^{-6}$ when $K_{\Lambda}<1.7$. The dashed lines show the two separatrixes obtained from the standard BKT flow equations derived from (9-13) to leading order in $K_{k}-K_{c}(\sigma)$ and $\sqrt{\lambda_{2,k}}$ [22]. The different colors of the trajectories are used to improve visibility and have no particular physical significance.
Short-range fixed point: Bose glass.
In addition to the BKT fixed point, the flow equations (9-13) admit a fully attractive short-range fixed point defined by $K^{*}=v^{*}=0$, where the long-range correlated part of the disorder is irrelevant: $\Delta_{2}^{*}(u)=0$ (Fig. 2, left panel). Since $\lim_{k\to 0}\theta_{k}=\theta^{*}=z^{*}-1$, which follows from a restoration of the statistical tilt symmetry at the fixed point, the superfluid stiffness $\rho_{s,k}=v_{k}K_{k}/\pi\sim k^{\theta^{*}+z^{*}-1}\to 0$ vanishes at the fixed point, but the compressibility $\kappa_{k}=K_{k}/\pi v_{k}\sim k^{\theta^{*}-z^{*}+1}\to{\rm const}$ remains finite. The system is not superfluid but compressible, as expected for a Bose glass.
The value of $\Delta_{1}^{*}(u)$ can be found analytically [24],
$\Delta_{1}^{*}(u)=\frac{1}{2\bar{L}_{0}}\left[\left(u-\frac{\pi}{2}\right)^{2}-\frac{\pi^{2}}{12}\right],$ (14)
and exhibits “cusps” at $u=0,\pm\pi,\pm 2\pi,\dots$ (Fig. 4). A cuspy functional form of the renormalized disorder correlator is known to be related to the existence of metastable states and glassy properties [37]. At nonzero momentum scale $k$, quantum tunneling between the ground state and these metastable states leads to a rounding of the cusp singularity into a quantum boundary layer. The latter controls the low-energy dynamics and is responsible for the $\sigma(\omega)\sim\omega^{2}$ behavior of the (dissipative) conductivity. We refer to Refs. [23, 24, 38] for a detailed discussion.
The stability of the short-range fixed point with respect to long-range disorder correlations can be determined by considering the linearized equation satisfied by $\Delta_{2,k}(u)$,
$k\partial_{k}\Delta_{2,k}(u)=(-3+\sigma)\Delta_{2,k}(u)-\bar{L}_{0}\Delta_{1}^{*}(0)\Delta^{\prime\prime}_{2,k}(u).$ (15)
Using $\Delta_{1}^{*}(0)=\pi^{2}/12\bar{L}_{0}$ , we obtain
$k\partial_{k}\lambda_{2,k}=\left(-3+\sigma+\frac{\pi^{2}}{3}\right)\lambda_{2,k}.$ (16)
The Bose-glass fixed point is therefore unstable when $\sigma<\sigma_{c}$, where $\sigma_{c}=3-\pi^{2}/3\simeq-0.289868$.
Long-range fixed point: Mott glass.
When $\sigma<\sigma_{c}$, we find that the localized phase is described by a fully attractive fixed point where $K^{*}=v^{*}=0$ and both disorder correlators $\Delta_{1}^{*}(u)$ and $\Delta_{2}^{*}(u)$ are nonzero (Fig. 2, right panel). In marked contrast to the short-range fixed point, $\theta^{*}\neq z^{*}-1$ (the statistical tilt symmetry is not restored at the fixed point), so that the compressibility $\kappa_{k}\sim k^{\theta^{*}-z^{*}+1}\to 0$ vanishes: The system is incompressible. Although the functions $\Delta_{1}^{*}(u)$ and $\Delta_{2}^{*}(u)$ cannot be found analytically, they can be obtained from the solution of the flow equations in the limit $k\to 0$ (Fig. 4). $\Delta_{1}^{*}(u)$ differs from (14) but still exhibits cusps at $u=n\pi$ ($n$ integer). This cusp singularity and the quantum boundary layer at nonzero RG scale $k$ describes the physics of rare low-energy metastable states and their coupling to the ground state by quantum fluctuations, as in the Bose-glass phase. Thus the only difference between the short-range and long-range fixed points is the vanishing of the compressibility in the latter case. The long-range fixed point describes an incompressible phase with a gapless conductivity $\sigma(\omega)\sim\omega^{2}$, i.e. a Mott glass.
Figure 4: Disorder correlators $\Delta_{1}^{*}(u)$ and $\Delta_{2}^{*}(u)$ at the short-range (SR) and long-range (LR) fixed points describing the Bose glass and the Mott glass, respectively.
The FRG approach fails for $\sigma\leq-0.5$ since the threshold function $\bar{L}_{x}$ diverges [22]. If we include the third cumulant $\Gamma_{3}$ in the three-replica-sum expansion (6), we find that some threshold functions diverge for $\sigma\to-1/3$. We conclude that the cumulant expansion of the functional $\Gamma_{k}[\{\phi_{a}\}]$ breaks down in the Mott glass. We argue below, considering an exactly solvable limit, that the existence of a Mott glass for all values $\sigma<\sigma_{c}$, as suggested by the FRG approach, is nevertheless likely to be correct.
The limit $\sigma=-1$ and $K\to 0^{+}$.
When $\sigma=-1$, the random potential $\xi(x)\equiv\xi=|\xi|e^{i\alpha}$ is perfectly correlated in both imaginary-time and space dimensions. For a given realization of the disorder, the action reads
$S=\int_{x,\tau}\left\{\frac{v}{2\pi K}\left[(\partial_{x}\varphi)^{2}+\frac{({\partial_{\tau}}\varphi)^{2}}{v^{2}}\right]-2|\xi|\cos(2\varphi)\right\}$ (17)
after a shift of $\varphi$ to eliminate the phase $\alpha$, and corresponds to a sine-Gordon model with a random coupling constant $|\xi|$. The BKT critical point for this model is located at $K_{c}=2$, which agrees with the result $K_{c}(\sigma)=3/2-\sigma/2$ obtained for an arbitrary $\sigma<0$. In the semiclassical limit $K\to 0^{+}$ [39], one can expand the cosine term about one of its minima, e.g. $\varphi=0$, to obtain a Gaussian action,
$S=\int_{x,\tau}\frac{1}{2\pi Kv}\bigl{[}v^{2}(\partial_{x}\varphi)^{2}+({\partial_{\tau}}\varphi)^{2}+m^{2}\varphi^{2}\bigr{]},$ (18)
where $m=(8\pi vK|\xi|)^{1/2}$ is a random mass. It is now straightforward to compute physical quantities and perform the disorder averaging with the Gaussian probability distribution defined by (3[22]. One finds that the compressibility vanishes whereas the conductivity is gapless. The ground state is a Mott glass as suggested by the FRG approach. Interestingly, for $\sigma=-1$ the conductivity vanishes as $|\omega|^{3}$ and not as $\omega^{2}$ as expected in a Bose glass and likely also in a Mott glass. Whether this is a peculiarity of the case $\sigma=-1$ —which could be explained by the fact that a disorder perfectly correlated in space is clearly not self-averaging, thus making the case $\sigma=-1$ special— or holds in the entire Mott glass phase is an open issue; another possibility would be a quantum phase transition at a critical value $-1<\sigma^{*}<\sigma_{c}$ between two Mott-glass phases with different low-frequency conductivities.
Although the first cumulant $W_{1}[J_{a}]=\overline{W[J_{a};\xi]}$ of the random functional $W[J;\xi]=\ln Z[J;\xi]$ is well defined —here $Z[J;\xi]$ is the partition function of the system and $J$ an external source that couples linearly to the field— we find that the second cumulant $W_{2}[J_{a},J_{b}]$ and its Legendre transform $\Gamma_{2}[\phi_{a},\phi_{b}]$ do not exist. We conclude that the cumulant expansion is not defined and there is even no guarantee that their generating functionals $\ln Z[\{J_{a}\}]$ and $\Gamma[\{\phi_{a}\}]$ exist. This is in line with the failure of the cumulant expansion observed in the FRG approach when $\sigma<\sigma_{c}$ and strongly suggests that the ground state is a Mott glass for all values of $\sigma<\sigma_{c}$.
Conclusion.
The Mott glass is a very intriguing phase of matter since it is associated with gapped single-particle excitations (which implies a vanishing compressibility) and gapless particle-hole excitations (hence the absence of gap in the conductivity). It has been pointed out that these two properties are compatible only in the presence of long-range interactions [40]. In the present study, long-range (elastic) interactions are a consequence of long-range disorder correlations.
There have been various proposals for the existence of a Mott glass in one-dimensional quantum fluids. It has been claimed that the interaction between disorder and a commensurable periodic potential can stabilize a Mott glass [41, 42], but this proposal does not take into account the necessary long-range interactions [40, 43]. It has also been claimed that the ground state of a disordered system with linearly confining interactions (disordered Schwinger model) is a Mott glass [44, 42] but this conclusion is in contradiction with another study [36]. The work reported in Ref. [45] predicts a Mott glass in a disordered Bose gas with long-range interactions; this conclusion however assumes a “jellium” model, where the global neutrality of the system is ensured by a uniform background of opposite charges, which is undoubtedly difficult to achieve experimentally.
Thus we believe that long-range disorder correlations appear as the most natural way proposed so far to observe a Mott glass in a one-dimensional Bose gas. Cold atoms are a natural platform for experimental study. Laser speckles allow to control, to some extent, the disorder potential acting on the gas [46, 47], but long-range power-law correlations seem out of reach. On the other hand, by generating the disorder potential from the spatial modulation of a laser beam using a digital micromirror device (DMD), the statistical properties of the disorder potential can be adjusted at will, which could open up the possibility of creating disorder potentials with the desired characteristics (see, e.g., Refs. [48, 49, 50]). Finally we point out that a classical analog of the Mott glass is expected to exist in heavy-ion-irradiated type-II superconductors with algebraically correlated columnar disorder [22].
We thank T. Chalopin, D. Clément and N. Cherroret for useful discussions on the experimental realization of random potentials with long-range correlations in ultracold gases, and E. Orignac for valuable comments.